Circular No. 7753

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2001 TU_{80} (LINEAR-NEAT) An object independently discovered and reported as nebulous with

a 3′′ coma by NEAT at Palomar on Nov. 16.5 UT has been identified by G. V. Williams, Minor Planet Center, with an apparently asteroidal object reported on Oct. 13 and 17 by LINEAR (m_{2} = 19.8; discovery observations on MPS 39482) and on Oct. 19 by NEAT at Haleakala (MPS 40840), from which an apparently routine minor-planet orbit was computed (MPEC 2001-V35). Following posting on the NEO Confirmation Page, P. Kušnirak at Ondřejov found the comet to be moderately condensed with a 0′.3 coma on CCD images taken on Nov. 17.2 with a 0.65-m f/3.6 reflector.

2001 UT α_{2000} δ_{2000} m_{1} Observer
Oct. 13.44483 7^{h}21^{m}41.^{s}06 +18^{o}15′03′′.1  LINEAR
Nov. 16.52608 8 14 55.59 +17 15 12.8 17.9 NEAT

Additional observations, together with the following orbital elements and an ephemeris, appear on MPEC 2001-W22:

T = 2001 Dec. 9.2409 TT  ω = 354.^{o}7750
e = 0.480203  Ω = 109.1832 }2000.0
q = 1.937019 AU  i = 6.5872
a = 3.726489 AU n^{o} = 0.1370106 P = 7.194 years

SUPERNOVA 2001fz IN NGC 2280 Y. L. Qiu and J. Y. Hu, Beijing Astronomical Observatory (BAO),

report the discovery by BAO Supernova Survey of an apparent supernova in unfiltered images taken with the BAO 0.6-m telescope on Nov. 15.8 (when the object was at mag ∼ 17.4) and 16.7 UT (mag ∼ 17.0). SN 2001fz is located at α = 6^{h}44^{m}50^{s}.26, δ = −27^{o}37′12′′.7 (equinox 2000.0), which is 15′′.8 east and 66′′.1 north of the nucleus of NGC 2280. An image taken with the same telescope on Nov. 8 showed no star at this position (limiting mag 18.0).

COMET P/2001 Q6 (NEAT) Total visual magnitude estimates: Oct. 23.08 UT, 11.0 (K. Sárneczky,

Ágasvár, Hungary, 0.44-m reflector); Nov. 5.75, 10.9 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 10.57, 10.8 (Y. Nagai, Yamanashi, Japan, 0.32-m reflector).

2001 November 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green